what is the equilibrium temperature when pouring 200 grams at 80 degrees Celsius while at room temperature into a 50 gram brass flask?
Answer #1

specific heat of brass is 380 J/Kg-K

specific heat of water is 4186J/Kg-K

Room temperature is 20oC

Heat gained by brass flask = Heat lost by water

MBCB(T-TB)=MwCw(Tw-T)

0.05*380*(T-20)=0.2*4186*(80-T)

856.2T=67356

T=78.67oC